OpenStudy (anonymous):

Jade measured the height of an art frame as 5.23 feet, but the actual height was 6 feet. What is the percent of error in Jade's measurement? 12.83% 13.55% 14.72% 15.55%

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8 @e.mccormick

OpenStudy (kamibug):

Subtract 5.23 from 6 then find what percent that is of 6. Here's a percent calc: http://www.onlineconversion.com/percentcalc.htm

OpenStudy (anonymous):

thx

jimthompson5910 (jim_thompson5910):

KamiBug is correct you use the rule described here http://www.mathsisfun.com/numbers/percentage-error.html

OpenStudy (anonymous):

@jim_thompson5910 is it A?

jimthompson5910 (jim_thompson5910):

yes, \[\Large p = \frac{|\text{approx} -\text{ exact}|}{\text{exact}} \times 100\%\] \[\Large p = \frac{|5.23-6|}{6} \times 100\%\] \[\Large p = \frac{|-0.77|}{6} \times 100\%\] \[\Large p = \frac{0.77}{6} \times 100\%\] \[\Large p \approx 0.128333 \times 100\%\] \[\Large p \approx 12.8333\%\]

OpenStudy (anonymous):

thank you!!!! :)

jimthompson5910 (jim_thompson5910):

np